OpenStudy (anonymous):

I have the answer can someone just see if I am correct? Rewrite the following function in the form f(x)=kx^p. f(x)=((2x^2)/(4x^4)) k= p=

OpenStudy (anonymous):

Could you please tell me your answers?

OpenStudy (anonymous):

k=-2 and p=-2

OpenStudy (anonymous):

Nope, that's incorrect. You got p correct, but k is incorrect.

OpenStudy (anonymous):

Is k=-8

OpenStudy (anonymous):

\[f(x)=\frac{2x^2}{4x^4}\] \[=\frac{x^2}{2x^4}\] \[=\frac{1}{2x^2}\] \[=\frac{1}{2}\times \frac{1}{x^2}\] \[=\frac{1}{2}\times x^{-2}\] \[=\frac{1}{2}x^{-2}\]

OpenStudy (anonymous):

Nope. Way off. Never just pull out a rabbit out of the hat and give me a random number.

OpenStudy (anonymous):

so then k=.5 and p=-2 sorry I typed in the to the calculator the wrong thing a dn just took the answer without looking at it first soorry and thank you

OpenStudy (anonymous):

No worries. And best you do these types without a calculator. That's something you should ponder about doing instead of using the calcualtor, because most calculators can't put x's and y's.